# Screened Thin-Target Bremsstrahlung with Partially-Ionized High-Z Species

### v1

@@ II.2 Sauter–Elwert Cross Section
$F_{\text{E}}\!\left(\bar{p}_{0},\bar{p},Z_{s}\right)=\frac{\xi}{\xi_{0}}\frac{1-\exp\!\left(-2\pi\xi_{0}\right)}{1-\exp\!\left(-2\pi\xi\right)},$ (17)
where $\xi_{0}=\alpha Z_{s}\bar{E}_{0}/\bar{p}_{0}=\alpha Z_{s}\beta_{0}^{-1}$ and $\xi=\alpha Z_{s}\beta^{-1}$. The resulting Sauter–Elwert (SE) cross section is then defined as
$\left(\frac{d^{2}\sigma_{e-i}^{Z_{s}}}{d\bar{k}\,d\Omega_{k}}\right)_{\!\text{SE}}^{\text{us}}:=F_{\text{E}}\!\left(\bar{p}_{0},\bar{p},Z_{s}\right)\left(\frac{d^{2}\sigma_{e-i}^{Z_{s}}}{d\bar{k}\,d\Omega_{k}}\right)_{\!\text{S}}^{\text{us}}.$ (18)
Physically, the Elwert factor $F_{\text{E}}$ accounts for the Coulomb distortion of the initial and final electron wave functions. It acts at the triply differential level where it modifies the local transition probability density. When $F_{\text{E}}$ is applied to the TDCS and one integrates over the outgoing electron angle $\Omega_{p}$ to obtain the DDCS, the resulting expression becomes essentially independent of the photon direction $\Omega_{k}$ in the neighborhood of the long-wavelength limit. This near-constancy allows $F_{\text{E}}$ to be treated as a global multiplicative factor that can be pulled out of the angular integration, yielding the compact form in Eq. (18). At shorter wavelengths, however, where the outgoing electron becomes slow and Coulomb distortion is strongest, this approximation breaks down and $F_{\text{E}}$ must in principle be retained inside the angular integrals. Elwert and Haug [7] addressed this by computing numerical Coulomb factors that extend the validity of the correction into the short-wavelength regime.
The analytic factor $F_{\text{E}}$ was obtained in the vicinity of the long-wavelength limit, where $\xi_{0}\!\simeq\!\xi$, and is nearly independent of the photon emission angle $\theta_{0}$ [24]. As $\xi_{0}$ decreases, this angular independence extends progressively toward the short-wavelength limit [8]. The Sauter–Elwert cross section is valid in the non-relativistic regime ($\xi_{0}\!\ll\!1$) and for sufficiently small $Z_{s}$. Unlike the Sauter result, it does not vanish at the tip region. For a more detailed discussion on Coulomb factors, we refer to [8, 7].
For highly relativistic incident electrons, and higher atomic numbers, the Sauter–Elwert corrected cross section becomes inadequate and exact numerical treatments of the Coulomb field, such as the Roche–Ducos–Proriol cross section, are required.
@@ II.3 Roche–Ducos–Proriol Cross Section
For high-$Z$ species ($Z_{s}>26$), the Sauter expression corrected within the first Born approximation (with Elwert factors) is no longer reliable, particularly in the tip region (see Sec. II.2). In such cases one must turn to formulations that treat the Coulomb field more precisely. The first analytic progress in this direction was made by Elwert and Haug, who derived their cross section using Furry–Sommerfeld–Maue (FSM) (usually simply called Sommerfeld–Maue (SM)) wave functions for the electron in the pure Coulomb potential [7]. This approach was later taken to the next order by Roche, Ducos, and Proriol, who added the dominant higher-order terms beyond Elwert–Haug [43]. The resulting RDP cross section remains the most accurate closed-form expression for bremsstrahlung in a pure Coulomb field in the energy range from a few to several tens of MeV, and it is the expression we employ for the first term on the right-hand side of Eq. (3). We do not delve in detail in the derivation of the result, since it is not within the scope of the present paper, and instead refer to the works from the original authors, as well as Furry [14] and Sommerfeld–Maue [48]. A good discussion of the former cross section can be found in the paper by Mangiarotti and Martins [32]. The RDP cross section reads
$\left(\frac{d^{3}\sigma_{e-i}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}\right)_{\text{RDP}}^{\text{us}}:=\frac{d^{3}\sigma_{\text{RDP},1}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}+\frac{d^{3}\sigma_{\text{RDP},2}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}.$ (19)
@@
$\frac{d^{3}\sigma_{\text{RDP},1}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}\equiv\frac{d^{3}\sigma_{\text{EH}}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}},$ (20)
@@
$-\left.\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)\right]\left|J_{1}\right|^{2}$
$+\left[\bar{E}_{0}\bar{E}+1+\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)\right]\left(\left|\bm{J}_{2}\right|^{2}+\left|\bm{J}_{3}\right|^{2}\right)$
@@
$+\left(\bm{J}_{2}\cdot\bar{\bm{p}}_{0}\right)\left(\bm{J}_{3}^{*}\cdot\bar{\bm{p}}\right)-\left(\bm{J}_{2}\cdot\bar{\bm{p}}\right)\left(\bm{J}_{3}^{*}\cdot\bar{\bm{p}}_{0}\right)$
@@
$+\left.\left.\left(\bm{J}_{2}\cdot\bm{J}_{3}^{*}\right)\left\{\bar{\bm{p}}_{0}\cdot\bar{\bm{p}}-\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)\right\}\right]\right\}$ (21)
and
$\frac{d^{3}\sigma_{\text{RDP},2}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}$ $=\alpha^{2}Z_{s}^{3}r_{e}^{2}\frac{\overline{k}\overline{p}\left(\bar{\bm{q}}\cdot\bar{\bm{k}}\right)}{\overline{p}_{0}\overline{q}D_{0}D}N$ (22)
@@
$+\left.\left.\bar{E}\left\{\bm{J}_{2}\cdot\bar{\bm{p}}_{0}-\left(\bm{J}_{3}\cdot\hat{\bm{k}}\right)\left(\bar{\bm{p}}_{\mathrm{0}}\cdot\hat{\bm{k}}\right)\right\}\right)e^{i\Phi}\right]$
$+\left.\alpha Z_{s}\frac{\bar{\bm{q}}\cdot\bar{\bm{k}}}{\overline{q}D_{0}D}\left[\bar{E}_{0}\bar{E}-1-\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)\right]\right\},$ (23)
where corresponding incident and outgoing electron’s momenta are $\bar{\bm{p}}_{0}=\bar{p}_{0}\bm{\hat{p}}_{0}$, $\bar{\bm{p}}=\bar{p}\bm{\hat{p}}$ respectively. The emitted photon momentum is $\bar{\bm{k}}=\bar{k}\bm{\hat{k}}$ and the recoil momentum of the target atom is $\bar{\bm{q}}=\bar{\bm{p}}_{0}-\bar{\bm{p}}-\bar{\bm{k}}$ as introduced in Sec. II.1. In Eqs. (21) and (23),
@@
where $a_{0}=\left(\bar{E}_{0}/\bar{p}_{0}\right)\alpha Z_{s}$ and $a=\left(\bar{E}/\bar{p}\right)\alpha Z_{s}$, while
$D_{0}$ $=2\left(\bar{E}_{0}\bar{k}-\bar{\bm{p}}_{0}\cdot\bar{\bm{k}}\right)\quad D=2\left(\bar{E}\bar{k}-\bar{\bm{p}}\cdot\bar{\bm{k}}\right)$ (25)
$\mu$ $=2\left(\bar{E}_{0}\bar{E}+\bar{p}_{0}\bar{p}-1\right)\quad\text{and}\quad x=1-\frac{D_{0}D}{\mu\bar{q}^{2}}.$ (26)
@@
$\Phi=a_{0}\ln\left(\frac{\bar{q}^{2}}{D}\right)-a\ln\left(\frac{\mu}{D}\right),$ (27)
@@
$J_{1}=$ $2\left(\frac{\bar{E}}{D_{0}}-\frac{\bar{E}_{0}}{D}\right)\frac{V+iaxW}{\bar{q}^{2}}$
$+2i\frac{\left(1-x\right)W}{D_{0}D}\left[\bar{E}_{0}a\left(\frac{\mu}{D}-1\right)-\bar{E}a_{0}\left(\frac{\mu}{D_{0}}+1\right)\right]$
$\bm{J}_{2}=$ $\frac{V+iaxW}{D\bar{q}^{2}}\bar{\bm{q}}-\frac{ia\left(1-x\right)W}{D_{0}D}\left[\left(\frac{\mu}{D}-1\right)\bar{\bm{q}}-\frac{\bar{\bm{P}}}{\bar{p}_{0}}\right]$
$\bm{J}_{3}=$ $\frac{V+iaxW}{D\bar{q}^{2}}\bar{\bm{q}}-\frac{ia_{0}\left(1-x\right)W}{D_{0}D}\left[\left(\frac{\mu}{D_{0}}+1\right)\bar{\bm{q}}-\frac{\bar{\bm{P}}}{\bar{p}}\right],$
with $V={}_{2}F_{1}\left(-ia_{0},ia;1;x\right)$ and $W={}_{2}F_{1}\left(1-ia_{0},1+ia;2;x\right)$. $_{2}F_{1}$ denotes the Gaussian hypergeometric function with argument $x\in[0,1]$ [2], while
$\bar{\bm{P}}=\bar{p}_{0}\bar{\bm{p}}+\bar{p}\bar{\bm{p}}_{\mathrm{0}}.$ (28)
Due to their high complexity, the expressions (21) and (23) can only be integrated over $\Omega_{p}$ numerically, and for that they may be reformulated as convenient sums:
$\frac{d^{3}\sigma_{\text{RDP},1}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}=\frac{d^{3}\sigma_{\text{EH}}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}=\frac{\alpha Z_{s}^{2}}{\pi^{2}}r_{e}^{2}\frac{\bar{k}\bar{p}}{\bar{p}_{0}}N\sum_{j=1}^{j=8}A_{j}^{\text{EH}},$ (29)
where
$A_{1}^{\text{EH}}=$ $\bar{E}_{0}\bar{E}-1-\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)J_{1}J_{1}^{*}$
$A_{2}^{\text{EH}}=$ $\left[\bar{E}_{0}\bar{E}+1+\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)\right]\left(\bm{J}_{2}\cdot\bm{J}_{2}^{*}+\bm{J}_{3}\cdot\bm{J}_{3}^{*}\right)$
[…]
and for the second term,
$\frac{d^{3}\sigma_{\text{RDP},2}^{Z_{s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}=\alpha^{2}Z_{s}^{3}r_{e}^{2}\frac{\overline{k}\overline{p}}{\overline{p}_{0}\overline{q}D_{0}D}N\left(\bar{\bm{q}}\cdot\bar{\bm{k}}\right)\sum_{j=1}^{j=2}A_{j}^{\text{RDP}},$ (30)
with
[…]
$A_{2}^{\text{RDP}}$ $=\alpha Z_{s}\frac{\bar{\bm{q}}\cdot\bar{\bm{k}}}{\overline{q}D_{0}D}\left[\bar{E}_{0}\bar{E}-1-\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)\right],$
and
$A_{11}^{\text{RDP}}$ $=\bar{E}_{0}\bar{E}-1-\left(\hat{\bm{k}}\cdot\bar{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot\bar{\bm{p}}\right)J_{1}$
[…]
@@
where $\vartheta_{pr}$ is the angle between $\bar{\bm{p}}$ and $\mathbf{r}$, and $r=\|\mathbf{r}\|$, the electron–nucleus separation [18]. This condition is satisfied for low atomic numbers at all angles and energies, but its validity is restricted to small angles for higher $Z$. In practice this approximation is sufficient since photon emission is concentrated in the forward direction at higher electron energies. The implications of this condition are further discussed in Sec. IV.
@@ III.1 Atomic Form Factor
From the Thomas–Fermi–Kirillov (TFK) or Yukawa (Y) approximate atomic potential models [26, 1] the form factor for the ion $Z_{s^{\prime},s}$, $F_{Z_{s^{\prime},s}}\left(\bar{q}\right)$, may be expressed in the simple form
$F_{Z_{s^{\prime},s}}\left(\bar{q}\right)=\frac{1}{Z_{s}}\frac{N_{s^{\prime},s}}{1+\left(\bar{q}\bar{a}_{Z_{s^{\prime},s}}/2\right)^{n}},$ (32)
where $n=3/2$ and $n=2$ for the TFK and Y potentials, respectively. Here $\overline{a}_{Z_{s^{\prime},s}}=2a_{Z_{s^{\prime},s}}/\alpha$ is the effective screening length, which depends on the atomic model and the ionization state. The parameter $N_{s^{\prime},s}$ denotes the number of bound electrons, so that the effective ionic charge participating in bremsstrahlung is $Z_{s}-N_{s^{\prime},s}=:Z_{s^{\prime},s}$. In the small-$\bar{q}$ limit, the form factor reduces to $F_{Z_{s^{\prime},s}}(\bar{q})\simeq N_{s^{\prime},s}$, in agreement with charge conservation.
@@
$F_{Z_{s^{\prime},s}}\left(\bar{q}\right)=\frac{N_{s^{\prime},s}}{Z_{s}}\sum_{i}\frac{\overline{A}_{s^{\prime},s,i}}{1+\left(\bar{q}\bar{a}_{Z_{s^{\prime},s,i}}/2\right)^{2}},$ (33)
where $\{A_{s^{\prime},s,i}\}_{i}$ denote the weights corresponding to each exponential for the different ionization levels. For normalization, we impose $\sum_{i}\overline{A}_{s^{\prime},s,i}=1$. This formulation provides a much more accurate description of screening effects, regardless of the ionization state or the atomic species. We refer to the work by Savoye-Peysson et al. for a more detailed description of the MY model [50]. Integration over $\phi$ and $\theta$ of the outgoing electron yields the DDCS in photon energy and emission angle
$\frac{d^{2}\sigma^{\text{MY}}_{Z_{s^{\prime},s}}}{d\bar{k}d\Omega_{k}}=\int_{0}^{\pi}\int_{0}^{2\pi}\frac{d^{3}\sigma^{\text{MY}}_{Z_{s^{\prime},s}}}{d\bar{k}d\Omega_{k}d\Omega_{p}}d\phi\sin\theta d\theta.$ (34)

### v2

@@ II.2 Sauter–Elwert Cross Section
$F_{\text{E}}\!\left({p}_{0},{p},Z\right)=\frac{\xi}{\xi_{0}}\frac{1-\exp\!\left(-2\pi\xi_{0}\right)}{1-\exp\!\left(-2\pi\xi\right)},$ (14)
where $\xi_{0}=\alpha Z{E}_{0}/{p}_{0}=\alpha Z\beta_{0}^{-1}$ and $\xi=\alpha Z\beta^{-1}$. The resulting Sauter–Elwert (SE) cross section is then defined as
$\left(\frac{d^{2}\sigma_{e-i}^{Z}}{d{k}\,d\Omega_{k}}\right)_{\!\text{SE}}^{\text{us}}:=F_{\text{E}}\!\left({p}_{0},{p},Z\right)\left(\frac{d^{2}\sigma_{e-i}^{Z}}{d{k}\,d\Omega_{k}}\right)_{\!\text{S}}^{\text{us}}.$ (15)
Physically, the Elwert factor $F_{\text{E}}$ accounts for the Coulomb distortion of the initial and final electron wave functions. It acts at the triply differential level where it modifies the local transition probability density. When $F_{\text{E}}$ is applied to the TDCS and one integrates over the outgoing electron angle $\Omega_{p}$ to obtain the DDCS, the resulting expression becomes essentially independent of the photon direction $\Omega_{k}$ in the neighborhood of the long-wavelength limit. As $\xi_{0}$ decreases, this angular independence extends progressively toward the short-wavelength limit [8]. This near-constancy allows $F_{\text{E}}$ to be treated as a global multiplicative factor that can be pulled out of the angular integration, yielding the compact form in Eq. (15). At shorter wavelengths, however, where the outgoing electron becomes slow and Coulomb distortion is strongest, this approximation breaks down and $F_{\text{E}}$ must in principle be retained inside the angular integrals. Elwert and Haug [7] addressed this by computing numerical Coulomb factors that extend the validity of the correction into the short-wavelength regime.
The Sauter–Elwert cross section is valid in the non-relativistic regime ($\xi_{0}\!\ll\!1$) and for sufficiently small atomic numbers $(Z\lesssim 26)$. Unlike the Sauter result, it does not vanish at the tip region. For a more detailed discussion on Coulomb factors, we refer to [8, 7].
For highly relativistic incident electrons, and higher atomic numbers, the Sauter–Elwert corrected cross section becomes inadequate and more advanced numerical treatments of the Coulomb field, such as the Roche–Ducos–Proriol cross section, are required.
@@ II.3 Roche–Ducos–Proriol Cross Section
For high-$Z$ species ($Z>26$), the Sauter expression corrected within the first Born approximation (with Elwert factors) is no longer reliable, particularly in the tip region (see Sec. II.2). In such cases one must turn to formulations that treat the Coulomb field more precisely. The first analytic progress in this direction was made by Elwert and Haug, who derived their cross section using Furry–Sommerfeld–Maue (FSM) (usually simply called Sommerfeld–Maue (SM)) wave functions [14, 50] for the electron in the pure Coulomb potential [7]. This approach was later taken to the next order by Roche, Ducos, and Proriol, who added the dominant higher-order terms beyond Elwert–Haug [45]. Despite its extensive use in the literature [18, 31, 32, 30], the RDP cross section as introduced in the original source [45] is not consistent, as its mixed-order character (combining third- and fourth-order contributions) does not necessarily lead to improved accuracy, as originally claimed. Although the mixed-order aspect of the initial formalism had been justified previously [18, 31, 32, 30], the resulting inconsistency was later identified and corrected at the level of the radiation matrix element by Jakubassa-Amundsen and Mangiarotti [20]. Here, we extend this correction to the level of the triply-differential cross section, yielding a consistently truncated third-order expression. The corrected RDP formalism provides the most accurate closed-form description of bremsstrahlung in a pure Coulomb field in the energy range from a few to several tens of MeV, and the integrated RDP TDCS is the expression we employ for the first term on the right-hand side of Eq. (3). The RDP cross section reads
$\left(\frac{d^{3}\sigma_{e-i}^{Z}}{d{k}d\Omega_{k}d\Omega_{p}}\right)_{\text{RDP}}^{\text{us}}:=\frac{d^{3}\sigma_{\text{RDP},1}^{Z}}{d{k}d\Omega_{k}d\Omega_{p}}+\frac{d^{3}\sigma_{\text{RDP},2}^{Z}}{d{k}d\Omega_{k}d\Omega_{p}}.$ (16)
@@
$\frac{d^{3}\sigma_{\text{RDP},1}^{Z}}{d{k}d\Omega_{k}d\Omega_{p}}\equiv\frac{d^{3}\sigma_{\text{EH}}^{Z}}{d{k}d\Omega_{k}d\Omega_{p}},$ (17)
@@
$-\left.\left(\hat{\bm{k}}\cdot{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot{\bm{p}}\right)\right]\left|J_{1}\right|^{2}$
$+\left[{E}_{0}{E}+1+\left(\hat{\bm{k}}\cdot{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot{\bm{p}}\right)\right]\left(\left|\bm{J}_{2}\right|^{2}+\left|\bm{J}_{3}\right|^{2}\right)$
@@
$+\left(\bm{J}_{2}\cdot{\bm{p}}_{0}\right)\left(\bm{J}_{3}^{*}\cdot{\bm{p}}\right)-\left(\bm{J}_{2}\cdot{\bm{p}}\right)\left(\bm{J}_{3}^{*}\cdot{\bm{p}}_{0}\right)$
@@
$+\left.\left.\left(\bm{J}_{2}\cdot\bm{J}_{3}^{*}\right)\left\{{\bm{p}}_{0}\cdot{\bm{p}}-\left(\hat{\bm{k}}\cdot{\bm{p}}_{0}\right)\left(\hat{\bm{k}}\cdot{\bm{p}}\right)\right\}\right]\right\},$ (18)
while the consistent third-order RDP term reads
$\frac{d^{3}\sigma_{\text{RDP},2}^{Z}}{d{k}d\Omega_{k}d\Omega_{p}}$ $=\alpha^{2}Z^{3}r_{e}^{2}\frac{{k}{p}\left({\bm{q}}\cdot{\bm{k}}\right)}{{p}_{0}{q}D_{0}D}N$ (19)
@@
$+\left.\left.{E}\left\{\bm{J}_{2}\cdot{\bm{p}}_{0}-\left(\bm{J}_{3}\cdot\hat{\bm{k}}\right)\left({\bm{p}}_{\mathrm{0}}\cdot\hat{\bm{k}}\right)\right\}\right)e^{i\Phi}\right].$ (20)
In Eqs. (18) and (20), $^{\bm{k}}={k}\bm{\hat{k}}$ and
@@
where $a_{0}=\left({E}_{0}/{p}_{0}\right)\alpha Z_{s}$ and $a=\left({E}/{p}\right)\alpha Z_{s}$, while
$D_{0}$ $=2\left({E}_{0}{k}-{\bm{p}}_{0}\cdot{\bm{k}}\right)\quad D=2\left({E}{k}-{\bm{p}}\cdot{\bm{k}}\right)$ (22)
$\mu$ $=2\left({E}_{0}{E}+{p}_{0}{p}-1\right)\quad\text{and}\quad x=1-\frac{D_{0}D}{\mu{q}^{2}}.$ (23)
@@
$\Phi=a_{0}\ln\left(\frac{{q}^{2}}{D}\right)-a\ln\left(\frac{\mu}{D}\right),$ (24)
@@
$J_{1}=$ $2\left(\frac{{E}}{D_{0}}-\frac{{E}_{0}}{D}\right)\frac{V+iaxW}{{q}^{2}}$
$+2i\frac{\left(1-x\right)W}{D_{0}D}\left[{E}_{0}a\left(\frac{\mu}{D}-1\right)-{E}a_{0}\left(\frac{\mu}{D_{0}}+1\right)\right]$
$\bm{J}_{2}=$ $\frac{V+iaxW}{D{q}^{2}}{\bm{q}}-\frac{ia\left(1-x\right)W}{D_{0}D}\left[\left(\frac{\mu}{D}-1\right){\bm{q}}-\frac{{\bm{P}}}{{p}_{0}}\right]$
$\bm{J}_{3}=$ $\frac{V+iaxW}{D{q}^{2}}{\bm{q}}-\frac{ia_{0}\left(1-x\right)W}{D_{0}D}\left[\left(\frac{\mu}{D_{0}}+1\right){\bm{q}}-\frac{{\bm{P}}}{{p}}\right].$
As for $V={}_{2}F_{1}\left(-ia_{0},ia;1;x\right)$ and $W={}_{2}F_{1}\left(1-ia_{0},1+ia;2;x\right)$, $_{2}F_{1}$ denotes the Gaussian hypergeometric function with argument $x\in[0,1]$ [1], while
${\bm{P}}={p}_{0}{\bm{p}}+{p}{\bm{p}}_{\mathrm{0}}.$ (25)
Due to their high complexity, the expressions (18) and (20) can only be integrated over $\Omega_{p}$ numerically. The evaluation of two hypergeometric functions at each point renders the computation demanding; for computational efficiency, the latter are implemented in the form of series expansions, and the results have been benchmarked against an arbitrary-precision library [23]. Details on the numerical implementation can be found in the source repository [16].
@@
where $\vartheta_{pr}$ is the angle between $\bm{p}$} and $\mathbf{r}$, and $r=\|\mathbf{r}\|$, the electron–nucleus separation [17]. This condition is satisfied for low atomic numbers at all angles and energies, but its validity is restricted to small angles for higher $Z$. In practice this approximation is sufficient since photon emission is concentrated in the forward direction at higher electron energies. The implications of this condition are further discussed in Sec. IV.
@@ III.1 Atomic Form Factor
From the Thomas–Fermi–Kirillov (TFK) or Yukawa (Y) approximate atomic potential models [25, 2] the form factor for the ion $Z_{s}$, $F_{Z_{s}}\left({q}\right)$, may be expressed in the simple form [^1] Note that the present definition of the form factor differs from Refs. [25, 48] by an overall factor of $1/Z$. Here we adopt the convention commonly used in the bremsstrahlung literature.
$F_{Z_{s}}\left({q}\right)=\frac{1}{Z}\frac{N_{s}}{1+\left({q}{a}_{Z_{s}}/2\right)^{n}},$ (27)
where $n=3/2$ and $n=2$ for the TFK and Y potentials, respectively. Here ${a}_{Z_{s}}=2\lambda_{Z_{s}}/\alpha$ is the effective screening length, with $\lambda_{Z_{s}}$ the screening parameter of the underlying Yukawa or TFK potential. It depends on the atomic model and the ionization state. In the small-q{q} limit, the form factor reduces to $F_{Z_{s}}({q})\simeq N_{s}/Z$, in agreement with charge conservation.
@@
$F_{Z_{s}}\left({q}\right)=\frac{N_{s}}{Z}\sum_{i}\frac{{A}_{s,i}}{1+\left({q}{a}_{Z_{s,i}}/2\right)^{2}},$ (28)
where $\{A_{s,i}\}_{i}$ denote the weights corresponding to each exponential for the different ionization levels. For normalization, we impose $\sum_{i}{A}_{s,i}=1$. This formulation provides an accurate description of screening effects, regardless of the ionization state or the atomic species. We refer to the work by Savoye-Peysson et al. for a detailed description of the MY model and its underlying assumptions [48]. Integration over $\phi$ and $\theta$ of the outgoing electron yields the DDCS in photon energy and emission angle
$\frac{d^{2}\sigma^{\text{MY}}_{Z_{s}}}{d{k}d\Omega_{k}}=\int_{0}^{\pi}\int_{0}^{2\pi}\frac{d^{3}\sigma^{\text{MY}}_{Z_{s}}}{d{k}d\Omega_{k}d\Omega_{p}}d\phi\sin\theta d\theta.$ (29)